# Quasi-randomization tests for network interference

### v2

@@ Abstract
Network interference amounts to the treatment status of one unit affecting the potential outcome of other units in the population. Testing for spillover effects in this setting makes the null hypothesis non-sharp. An interesting approach to tackling the non-sharp nature of the null hypothesis in this setup is constructing conditional randomization tests such that the null is sharp on the restricted population. Such approaches can pose computational challenges as finding these appropriate sub-populations based on experimental design can involve solving an NP-hard problem. In this paper, we view the network amongst the population as a random variable instead of being fixed. We propose a new approach that builds a conditional quasi-randomization test. We build the (non-sharp) null distribution of no spillover effects using random graph null models. We show that our method is exactly valid in finite samples under mild assumptions. Our method displays enhanced power over other methods, substantially improving cluster randomized trials. We illustrate our methodology to test for interference in a weather insurance adoption experiment run in rural China.
Keywords: Causal Inference, Network Interference, Peer Effects, Randomization Tests, Random Graphs
@@ 1 Introduction
A common assumption in causal inference is that of the Stable Unit Treatment Value Assumption (SUTVA). SUTVA assumes no dependence of a unit’s outcome on another unit’s treatment status. While this allows for unbiased treatment effect estimation and valid inferential strategies, there are many experimental and observational settings where interference between units takes place (Cox (1958)). There are multiple pathways via which interference can take place in the population. It can result from the spread of treatment amongst the population units, the respective outcomes of treated units affecting the outcomes of others, or a combination of both. For example, the spread of a contagious disease depends on the number of infected people in the population (Halloran and Struchiner (1995)). Thus, the estimation of any preventive measures taken will be challenged because of a likely SUTVA violation. Consider a population-level policy evaluation program, like an insurance product campaign (Cai et al. (2015)). Due to social network ties in the population, information dissemination is a plausible interference mechanism that can occur and affect evaluated estimates. Similar phenomena can be seen amongst experiments run on social media networks, like Linkedin, or online marketplaces like eBay, owing to the interconnected nature of these platforms (Pouget-Abadie et al. (2019); Blake and Coey (2014)). The total treatment effect, in such cases, can be decomposed into two effects, one of the direct effects and the other of spillover effects (Sobel (2006)).
@@
The classical method for testing treatment effects is the “Fisher Randomization Test” or FRT (Fisher (1936)). FRT is used to test for a sharp null hypothesis of no treatment effect for all units in the population. The test exploits the imputability of counterfactuals under the null. Imputation enables the determination of the randomization null distribution of a test statistic of choice. If one were to test for spillover effects, it would render the null non-sharp, leading to imputability issues for the test statistic. There is a growing body of literature that addresses this issue by considering a carefully determined subset of units in the population and treatment assignment vector, such that on these selected units and treatment vectors, the null is sharp (Aronow and Samii (2017); Athey et al. (2018); Basse et al. (2019b)). Such conditional randomization tests also lead to valid exact p-values, as with the Fisher Randomization Test. Basse et al. (2019b) build a permutation test under the particular case of clustered interference. While a permutation test is computationally cheaper, the method lacks generalizability because of the structured nature of the interference considered. Aronow and Samii (2017) and Athey et al. (2018) generalize the above for a given network and evaluate a range of non-sharp null hypotheses for testing. Puelz et al. (2022) presents a graph-theoretic approach to constructing the conditioning mechanism of a sub-population of units and the corresponding treatment assignment vectors for which the null is sharp. However, the proposed methodologies lose power in cluster randomized experiments.
@@
This paper considers the network to be coming from a distribution itself. This idea contrasts with the setup considered in the current methods, assuming the network is fixed. We draw a parallel here to estimating the probability of unit treatment assignment via propensity scores in the case of observational studies. The propensity of two nodes in a network being connected can be considered a function of their network characteristics, covariates, or both. This leads us to think of these linkages as coming from a generating process with an associated probability distribution. The current methods are based on strict randomization inference with no distributional assumptions for either the network or the potential outcomes, achieving a generalization of the Fisher Randomization Test. This allows for the widespread validity of the methods proposed, albeit it might be at the cost of power. This paper explores how leveraging appropriate stochastic assumptions can significantly boost power and lead to methodological advancements.
@@
We assume that a given vector of sufficient statistics characterizes the network data-generating process and can thus be represented as an exponential random graph model. We first take the degree sequence of the graph as a sufficient statistic. Proposed in Newman et al. (2001), random graphs with arbitrary degree distribution represent real-world social networks closely owing to their ability to capture heterogeneity in degree. This assists us in building a quasi-randomization test based on the network distribution by randomizing over graphs from this generating process. We show that conditional on the observed treatment, a test statistic’s (quasi) randomization distribution leads to valid p-values. We then generalize the model to a vector of sufficient statistics and present a method that works for any set of such network statistics.
@@
Model-based approaches with distribution assumptions on outcome model function have been previously considered in the literature for identifying and estimating peer effects (Manski (1993, 2013); Bowers et al. (2013); Toulis and Kao (2013); Blume et al. (2015)). More recently, Li and Wager (2022) and Leung (2020) proposed estimation strategies for spillover effects under a distributional assumption on the network formation model. The inferential methods for the proposed estimation procedures involve non-standard or asymptotic approximations (Imbens and Manski (2004)). With a careful identification of causal estimates, such large sample normal approximations of confidence interval conform to a Neymanian perspective.
@@
Our first contribution is to extend the random graph interference model into the potential outcome framework. We generalize and define sharp/non-sharp null hypotheses for random graph models. Building upon this, we develop a randomization-based test for a given random graph model. This leads to finite sample validity of the p-value obtained, in line with Fisherian procedures.
@@
As a second contribution, our method addresses several technical challenges encountered in the current methods. We show that randomization tests based on network-generating processes lead to imputable test statistics under the null, conditional on the treatment assignment. This overcomes a significant challenge in randomization methods based on random treatment assignment. The state-of-the-art techniques restrict attention to a carefully chosen subset of focal units and claim null of no spillover effects for only these units. In contrast, our method tests for a stricter class of nulls with no spillover effects on all the units.
@@
Another hurdle the current methodologies face is the loss of power in cluster randomized trials. Determination of conditional randomization distribution given a conditioning mechanism is not straightforward. The restriction of fixed focal units in the randomization null reduces the sample space of possible treatment assignment vectors in the null. It can lead to an enormous loss of power or render the null degenerate completely. A burgeoning literature set focuses on improving current methods by incorporating design-based techniques (Puelz et al. (2022); Basse et al. (2019a)). Unlike distribution-free methods, we show how incorporating information about network generation into the methodology can boost power.
@@
Our third significant contribution is the computational efficiency of the proposed method. We suggest that sufficient network statistics can characterize the random graph model. We first take a single sufficient statistic as the network’s degree sequence and build a permutation test by generating the null over the set of graphs with the same degree sequence as the observed graph. Random graph generation of a prescribed degree sequence admits fast and efficient polynomial time algorithms for large graphs (Viger and Latapy (2005); Fosdick et al. (2018)). Thus, we obtain valid randomization tests via permutation of edges of the graph in a manner that preserves degree sequence. We then present a method for a general exponential random graph model characterized by an arbitrary set of network statistics and build a permutation test by generating graphs in the same isomorphism class under the null. This is closely related to the literature on the random graph null model used in network analysis with applications in protein structure discovery and community detection among many (Milenković et al. (2009); Sah et al. (2014)).
@@
The remaining paper is structured as follows. In Section 2, we set up notation and formalize the null hypothesis of interest, incorporating network random variable. Section 3 presents a randomization test based on a given network-generating process and its application to different experimental designs. Section 4 extends this further to random graph models and presents our central methodology introducing the permutation test. In Section 5, we present extensive numerical simulations to evaluate the performance of the proposed test. Sections 6 and 7 illustrate our methodology as an empirical application and conclude the paper. Proofs, algorithms, and additional numerical results are presented in the Appendix.
@@ 2.1 Notation
Consider a population =$\mathbb{P}=\{1,2,...,N\}$ where $N$ is the total number of units in the population. Consider a vector $Z\in{\mathbb{Z}}^{N}$ where each element in the vector indicates treatment assigned to the corresponding indexed unit in the population. Here, $\mathbb{Z}$ represents the set of possible treatments to which the population units can be exposed. We take =$\mathbb{Z}=\{0,1\}$ for this article. Hence, any treatment assignment vector is an N-dimensional binary vector. We consider the population $\mathbb{P}$ to be connected via an undirected graph $G=(\mathbb{P},E)$ where $E\subseteq\mathbb{P}\times\mathbb{P}$. For distinct units $i$ and $j$ such that $(i,j)\in E$ represents a connection between the two units. With some slight abuse of notation, we also represent the graph’s adjacency matrix by $G$. Thus, $G$ is an $N\times N$ binary matrix. Here, for some $i$ and $j$, $G_{ii}=0$ and $G_{ij}=G_{ji}$ owing to $G$ being a undirected graph with no self-loops. If $G_{ij}=1$, then we say unit $i$ and $j$ are neighbours. This can also be denoted by $i\in\mathcal{N}_{G}(j)$ or vice-versa. Here, $\mathcal{N}_{G}(i)$ represents all the neighbors of unit $i$ in graph $G$. We may write $\mathcal{N}(i)$ if the graph is clear in context. We define the total number of unit neighbors $i$ as the degree of $i$. This is denoted by $deg(i)$ and equals $|N(i)|$. We denote the distance between two units $i$ and $j$ as $dist(i,j)$. It is defined as the shortest path length between the two units. By convention, we take $dist(i,i)=0$ and $dist(i,j)=\infty$ if no path exists between the two distinct units. Consider $\mathbb{P^{\prime}}\subseteq\mathbb{P}$. We denote $G[\mathbb{P^{\prime}}]$ as the sub-graph induced by $\mathbb{P^{\prime}}$. That is, $G[\mathbb{P^{\prime}}]=(\mathbb{P^{\prime}},E^{\prime})$ where $E^{\prime}=\{(i,j)\in\mathbb{P^{\prime}}\times\mathbb{P^{\prime}}|\>(i,j)\in E\}$.
@@
Each unit $i$ in the population is ‘characterized’ by a set of pre-treatment or covariate vectors given by $X_{i}$ where $X_{i}$ is a m-dimensional vector. $X_{i}$ is fixed and influences the outcome via the network’s potential outcome function or formation. Let $X$ denote the $N\times M$ matrix where each row corresponds to the respective unit’s covariate vector. Consider the treatment vector to be drawn from a probability distribution $P_{X}(Z,G):\{0,1\}^{N}\times\{0,1\}^{N\times N}\rightarrow[0,1]$. We define the conditional probability distribution $P(Z|G):\{0,1\}^{N}\rightarrow[0,1]\subset\mathbb{R}$ as $P_{Z|G}$. Let $Z_{obs}$ be the realized treatment drawn from the treatment assignment distribution $P_{Z|G}$. Consider the graph $G$ to be coming from the marginal probability distribution $P(G):\{0,1\}^{N\times N}\rightarrow[0,1]\subset\mathbb{R}$, denoted as $P_{G}$. We denote the observed sample from the distribution $P_{G}$ to be $G_{obs}$. We define the potential outcome function $Y:\{0,1\}^{N}\times\{0,1\}^{N\times N}\rightarrow\mathbb{R}^{N}$ as a real-valued function that takes treatment assignment and network adjacency matrix as arguments. We will consider $Y$ to be fixed, and each element $i$ of an image in $Y(Z,G)$, $Y_{i}(Z,G)$ denotes the outcome of $i^{th}$ unit in the population under the treatment assignment $Z$ and network assignment $G$. We denote the corresponding potential outcome as $Y_{obs}$ that is, $Y_{obs}=Y(Z_{obs},G_{obs})$. We note that the potential outcome function implicitly contains information in $X$ and can be written as $Y_{X}(Z,G)$. For the remainder of the article, we will use the notation $Y(Z,G)$ for readability.
@@ 2.2 Definitions
The assignment mechanism defines the joint probability distribution of treatment assignment vectors given covariates and potential outcomes, denoted by $P(Z|X,Y(0),Y(1))$. The following are three key assumptions that are considered in the causal inference literature for assignment mechanism: (i) individualistic assignment, which restricts the treatment assignment of a unit to its covariates and outcomes (ii) probabilistic assignment, which assumes non-integer probabilities of treatment assignment to units (iii) unconfoundedness assumption, which restricts the dependence of the probability assignment of units to only covariates, and not the potential outcome (Imbens and Rubin (2015)). Given a well-defined potential outcome function, individualistic assignment mechanism assumption constitutes Stable Unit Treatment Value Assumption (SUTVA). We argue that the underlying network structure among the population is an implicit form of treatment. Hence, we generalize the treatment to a bivariate treatment assignment vector $(Z,G)$. We assume that the joint treatment assignment is unconfounded; that is,
$P(Z,G|X,Y(0),Y(1))=P(Z,G|X).$
@@
[…]

### v3

@@ Abstract
Network interference amounts to the treatment status of one unit affecting the potential outcome of other units in the population. Testing for spillover effects in this setting makes the null hypothesis non-sharp. An interesting approach to tackling the non-sharp nature of the null hypothesis in this setup is constructing conditional randomization tests such that the null is sharp on the restricted population. In randomized experiments, conditional randomized tests hold finite sample validity and are assumption-lean. In this paper, we incorporate the network amongst the population as a random variable instead of being fixed. We propose a new approach that builds a conditional quasi-randomization test. To build the (non-sharp) null distribution of no spillover effects, we use random graph null models. We show that our method is exactly valid in finite samples under mild assumptions. Our method displays enhanced power over state-of-the-art methods, with a substantial improvement in cluster randomized trials. We illustrate our methodology to test for interference in a weather insurance adoption experiment run in rural China.
Keywords: Causal Inference, Network Interference, Peer Effects, Randomization Tests, Random Graphs.
@@ 1 Introduction
A common assumption in causal inference is that of the Stable Unit Treatment Value Assumption (SUTVA). SUTVA assumes no dependence of a unit’s outcome on another unit’s treatment status. While this allows for unbiased treatment effect estimation and valid inferential strategies, there are many experimental and observational settings where interference between units takes place (Cox (1958)). There are multiple pathways via which interference can take place in the population. It can result from the spread of treatment amongst the population units, the respective outcomes of treated units affecting the outcomes of others, or a combination of both. For example, the spread of a contagious disease depends on the number of infected people in the population (Halloran and Struchiner (1995)). Thus, the estimation of any preventive measures taken will be challenged because of a likely SUTVA violation. Consider a population-level policy evaluation program, like an insurance product campaign (Cai et al. (2015)). Due to social network ties in the population, information dissemination is a plausible interference mechanism that can occur and affect the evaluated estimates. Similar phenomena can be seen amongst experiments run on social media networks, like LinkedIn, or online marketplaces like eBay, owing to the interconnected nature of these platforms (Pouget-Abadie et al. (2019); Blake and Coey (2014)). The total treatment effect, in such cases, can be decomposed into two effects, one of the direct effects and the other of spillover effects (Sobel (2006)).
@@
The classical method for testing treatment effects is the “Fisher Randomization Test” or FRT (Fisher (1936)). FRT is used to test for a sharp null hypothesis of no treatment effect for all units in the population. The test exploits the imputability of counterfactuals under the null. Imputation enables the determination of the randomization null distribution of a test statistic of choice. If one were to test for spillover effects, it would render the null non-sharp, leading to imputability issues for the test statistic. There is a growing body of literature that addresses this issue by considering a fixed subset of units in the population and treatment assignment vector, called ‘focal units’, such that on these focal units the null is sharp (Aronow and Samii (2017); Athey et al. (2018); Basse et al. (2019b)). Such conditional randomization tests also lead to valid exact p-values, as with the Fisher Randomization Test. Basse et al. (2019b) builds a permutation test under clustered interference. Aronow and Samii (2017) and Athey et al. (2018) generalize the above for a given network and evaluate a range of non-sharp null hypotheses for testing. Puelz et al. (2022) presents a graph-theoretic approach to constructing the conditioning mechanism of selecting the focal units and the corresponding treatment assignment vectors for which the null is sharp. While the above considers the problem of testing for interference in an experimental setting, the literature on testing for spillover effects in an observational setting is still in its infancy.
@@
This paper considers the network to be coming from a distribution itself. This idea expands upon the setup considered in the current methods, where the network is taken to be fixed. We draw a parallel here to estimating the probability of unit treatment assignment via propensity scores in the case of observational studies. The propensity of two nodes in a network being connected can be considered a function of their network characteristics, covariates, or both. This leads us to think of these linkages as coming from a generating process with an associated probability distribution. The current methods are based on strict randomization inference with no distributional assumptions for either the network or the potential outcomes, achieving a generalization of the Fisher Randomization Test. This allows for the widespread validity of the methods proposed. Since the experimenter cannot randomize this network, we explore how to incorporate appropriate stochastic assumptions on the network in the context of a conditional randomization test.
@@
We assume that a given vector of sufficient statistics characterizes the network data-generating process and can thus be represented as an exponential random graph model. We first take the degree sequence of the graph as a sufficient statistic. Proposed in Newman et al. (2001), random graphs with arbitrary degree distribution represent real-world social networks closely owing to their ability to capture heterogeneity in degree. This assists us in building a quasi-randomization test based on the network distribution by randomizing over graphs from this generating process. We show that, conditional on the observed treatment, a test statistic’s (quasi) randomization distribution leads to valid p-values. We then generalize the model to a vector of sufficient statistics and present a method that works for any set of such network statistics.
@@
Model-based approaches with distribution assumptions on outcome model function have been previously considered in the literature for identifying and estimating peer effects (Manski (1993, 2013); Bowers et al. (2013); Toulis and Kao (2013); Blume et al. (2015)). More recently, Li and Wager (2022) and Leung (2020) proposed estimation strategies for spillover effects under a distributional assumption on the network formation model. The inferential methods for the proposed estimation procedures involve asymptotic approximations (Imbens and Manski (2004)). Such large sample normal approximations of confidence interval conform to a Neymanian perspective.
@@
In line with Fisherian procedures, our first contribution is to extend the random graph interference model into the potential outcome framework. We generalize and define sharp/non-sharp null hypotheses for random graph models. Building upon this, we develop a randomization-based test for a given random graph model. This leads to finite sample validity of the p-value obtained.
@@
As a second contribution, we show that randomization tests based on network-generating processes lead to imputable test statistics under the null, conditional on the treatment assignment and network covariates. This overcomes a technical challenge in randomization methods based on random treatment assignment, where the null of no spillover effects is tested only upon the focal units (Athey et al. (2018)). In contrast, our method tests for a stricter class of nulls with no spillover effects on the given population.
@@
Another hurdle the current methodologies face is low power. Determination of the conditional randomization distribution given a conditioning mechanism is not straightforward. The restriction of fixed focal units in the randomization null reduces the sample space of possible treatment assignment vectors in the null. It can lead to an enormous loss of power or render the null degenerate completely. A burgeoning literature set focuses on improving current methods by incorporating design-based techniques for selecting conditioning mechanisms (Puelz et al. (2022); Basse et al. (2019a)). We show how incorporating information about network generation into the methodology can aid in power.
@@
Our methods are also computationally feasible. We suggest that sufficient network statistics can characterize the random graph model. We first take a single sufficient statistic as the network’s degree sequence and build a permutation test by generating the null over the set of graphs with the same degree sequence as the observed graph. Random graph generation of a prescribed degree sequence admits fast and efficient polynomial time algorithms for large graphs (Viger and Latapy (2005); Fosdick et al. (2018)). Thus, we obtain valid randomization tests via the permutation of edges of the graph in a manner that preserves the degree sequence. We then present a method for a general exponential random graph model characterized by an arbitrary set of network statistics and build a permutation test by generating degree-preserving graphs in the same isomorphism class under the null. This is closely related to the literature on the random graph null model used in network analysis with applications in protein structure discovery and community detection among many (Milenkovic et al. (2009); Sah et al. (2014)).
@@ 2.1 Notation
Consider a population $\mathbb{P}=\{1,2,...,N\}$ where $N$ is the total number of units in the population. Consider a vector $Z\in{\mathbb{Z}}^{N}$ where each element in the vector indicates treatment assigned to the corresponding indexed unit in the population. Here, $\mathbb{Z}$ represents the set of possible treatments to which the population units can be exposed. We take $\mathbb{Z}=\{0,1\}$ for this article. Hence, any treatment assignment vector is an N-dimensional binary vector. We consider the population $\mathbb{P}$ to be connected via an undirected graph $G=(\mathbb{P},E)$ where $E\subseteq\mathbb{P}\times\mathbb{P}$. For distinct units $i$ and $j$ such that $(i,j)\in E$ represents a connection between the two units. With some slight abuse of notation, we also represent the graph’s adjacency matrix by $G$. Thus, $G$ is an $N\times N$ binary matrix. Here, for some $i$ and $j$, $G_{ii}=0$ and $G_{ij}=G_{ji}$ owing to $G$ being a undirected graph with no self-loops. If $G_{ij}=1$, then we say unit $i$ and $j$ are neighbours. This can also be denoted by $i\in\mathcal{N}_{G}(j)$ or vice-versa. Here, $\mathcal{N}_{G}(i)$ represents all the neighbors of unit $i$ in graph $G$. We may write $\mathcal{N}(i)$ if the graph is clear in context. We define the total number of unit neighbors $i$ as the degree of $i$. This is denoted by $deg(i)$ and equals $|N(i)|$. We denote the distance between two units $i$ and $j$ as $dist(i,j)$. It is defined as the shortest path length between the two units. By convention, we take $dist(i,i)=0$ and $dist(i,j)=\infty$ if no path exists between the two distinct units. Consider $\mathbb{P^{\prime}}\subseteq\mathbb{P}$. We denote $G[\mathbb{P^{\prime}}]$ as the sub-graph induced by $\mathbb{P^{\prime}}$. That is, $G[\mathbb{P^{\prime}}]=(\mathbb{P^{\prime}},E^{\prime})$ where $E^{\prime}=\{(i,j)\in\mathbb{P^{\prime}}\times\mathbb{P^{\prime}}|\>(i,j)\in E\}$.
@@
Each unit $i$ in the population is ‘characterized’ by a set of pre-treatment or covariate vectors given by $X_{i}$, where $X_{i}$ is an m-dimensional vector. $X_{i}$ is fixed and influences the outcome via the network’s potential outcome function or formation. Let $X$ denote the $N\times M$ matrix where each row corresponds to the respective unit’s covariate vector. Consider the treatment vector to be drawn from a probability distribution $P_{X}(Z,G):\{0,1\}^{N}\times\{0,1\}^{N\times N}\rightarrow[0,1]$. We define the conditional probability distribution $P(Z|G):\{0,1\}^{N}\rightarrow[0,1]\subset\mathbb{R}$ as $P_{Z|G}$. Let $Z_{obs}$ be the realized treatment drawn from the treatment assignment distribution $P_{Z|G}$. Consider the graph $G$ to be coming from the marginal probability distribution $P(G):\{0,1\}^{N\times N}\rightarrow[0,1]\subset\mathbb{R}$, denoted as $P_{G}$. We denote the observed sample from the distribution $P_{G}$ to be $G_{obs}$. We define the potential outcome function $Y:\{0,1\}^{N}\times\{0,1\}^{N\times N}\rightarrow\mathbb{R}^{N}$ as a real-valued function that takes treatment assignment and network adjacency matrix as arguments. We will consider $Y$ to be fixed, and each element $i$ of an image in $Y(Z,G)$, $Y_{i}(Z,G)$, denotes the outcome of the $i^{th}$ unit in the population under the treatment assignment $Z$ and network assignment $G$. We denote the corresponding potential outcome as $Y_{obs}$ that is, $Y_{obs}=Y(Z_{obs},G_{obs})$. We note that the potential outcome function implicitly contains information in $X$ and can be written as $Y_{X}(Z,G)$. For the remainder of the article, we will use the notation $Y(Z,G)$ for readability unless deemed necessary.
@@ 2.2 Definitions
The assignment mechanism defines the joint probability distribution of treatment assignment vectors given covariates and potential outcomes, denoted by $P(Z|X,Y(0),Y(1))$. The following are three key assumptions that are considered in the causal inference literature for assignment mechanism: (i) individualistic assignment, which restricts the treatment assignment of a unit to its covariates and outcomes (ii) probabilistic assignment, which assumes non-integer probabilities of treatment assignment to units (iii) unconfoundedness assumption, which restricts the dependence of the probability assignment of units to only covariates, and not the potential outcome (Imbens and Rubin (2015)). Given a well-defined potential outcome function, the individualistic assignment mechanism assumption constitutes Stable Unit Treatment Value Assumption (SUTVA). We argue that the underlying network structure among the population is an implicit form of treatment. Hence, we generalize the treatment to a bivariate treatment assignment vector $(Z,G)$. We assume that the joint treatment assignment is unconfounded; that is,
$P(Z,G|X,Y(0),Y(1))=P(Z,G|X).$ (1)
Remark 1.
It should be noted that the graph $G$ only maps the interference structure in the population. Any other dependence of the potential outcome function on the network characteristics and other confounding factors comprises the covariates in the study. We denote the covariates corresponding to network statistics as $G_{X}$, which may be dependent on G. Apart from $G_{X}$, in this article, we restrict the potential outcome function to be covariate-free for the sake of brevity. Thus, Equation 1 can be written as
$P(Z,G|X,Y(0),Y(1))$ $=P(Z,G|G_{X},Y(0),Y(1)),$
$=P(Z,G|G_{X}).$
For further details on covariate balancing in randomized experiments, refer to Liu and Yang (2020).
@@
$\text{s.t.}\quad Z\cdot D_{G}^{(k-1)}(i)=Z^{\prime}\cdot D_{G^{\prime}}^{(k-1)}(i).$